OpenStudy (anonymous):

Find the least common multiple of x^3 - x^2 + x - 1 and x^2 - 1. Write the answer in factored form

OpenStudy (anonymous):

is it (x+1)(x-1)(x^2 - 1)

OpenStudy (lgbasallote):

isnt (x+1)(x-1) = (x^2 - 1)?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so would it be (x^2 + 1)

OpenStudy (lgbasallote):

first step factor x^2 - x^2 + x - 1 = (x-1)(x^2 + 1) x^2 - 1 = (x+1)(x-1) therefore LCM would be (x+1)(x-1)(x^2 + 1) so yeah it's (x^2 + 1)(x^2 - 1) but you can simplify that if you want hehe

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

welcome :D